OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24 I have this so far but I don't know how to get the solution. 2- 3x-6y=12 -3x -3x y= -3x+12 ---------- -6 9x+2y= -24 -9x -9x y= -9x-24 --------- 2

OpenStudy (fibonaccichick666):

Do you understand how to graph and get the answer?

OpenStudy (anonymous):

No I suck at this :(

OpenStudy (fibonaccichick666):

Don't say that. Ok so do you know how to graph the first equation?

OpenStudy (anonymous):

I'm guessing I would graph it at 0,-3?

OpenStudy (anonymous):

or (0,-6)

OpenStudy (fibonaccichick666):

Let's start by solving for y. That will make this easier :) (I know you did it already but I'd like you to do it again) Please show all of your steps no matter how insignificant they may seem.

OpenStudy (anonymous):

3x – 6y = 12 -3x -3x -6y= -3x+12 ---- ------ -6 -6 y= -3x +12 ------- -6 9x + 2y = –24 -9x -9x 2y= -9x - 24 --- ------- 2 2 y= -9x - 24 ----- 2

OpenStudy (fibonaccichick666):

ok so I have a quick demo if you like http://www.educreations.com/lesson/view/jaylene/9305370/

OpenStudy (fibonaccichick666):

But that looks good. an you distribute the -6 for the first eq. then simplify?

OpenStudy (anonymous):

y= -3x - 6 --- -6 y= -9x -12 --- 2

OpenStudy (fibonaccichick666):

not quite. What is \(\frac{12}{6}=?\)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so its y = -3x - 2 ---- -6

OpenStudy (fibonaccichick666):

so the first eq. we have \[y=\frac{-3x+12}{-6}\] We distribute the -6 to get: \[y=\frac{-3x}{-6} +\frac{12}{-6}\]

OpenStudy (fibonaccichick666):

Can you simplify this?

OpenStudy (anonymous):

y =-3y - 2 ----- -6

OpenStudy (fibonaccichick666):

simplify the x term please

OpenStudy (anonymous):

y = 2x - 2

OpenStudy (fibonaccichick666):

\(\frac{3}{6}=2?\)

OpenStudy (anonymous):

0.5 but its a decimal?

OpenStudy (fibonaccichick666):

yep, or one half either representation works

OpenStudy (fibonaccichick666):

so when x=0 y=?

OpenStudy (anonymous):

-2? idk

OpenStudy (fibonaccichick666):

that's right good guess

OpenStudy (fibonaccichick666):

just plug in x=0 then whatever you get is your answer.

OpenStudy (fibonaccichick666):

that is how you find points to plot for a graph

OpenStudy (anonymous):

but how do i get the solution of both equations?

OpenStudy (fibonaccichick666):

You will do the same thing with the other eq. solve for y (also referenced as the standard form of a line) then graph. After you graph, you will notice that the lines intersect. On a graph, that is your answer. The coordinates of that point. It is where the lines are equal.

OpenStudy (anonymous):

i got (4,1)

OpenStudy (fibonaccichick666):

I do not agree. I got x=-5

OpenStudy (anonymous):

so its (-5,1) ?

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

did you graph it?

OpenStudy (anonymous):

yeah but since the second equation is y=- 9 x - 12 made it hard so i can't see it clearly --- 2

OpenStudy (fibonaccichick666):

Yes the second eq is \(y=\frac{-9}{2}x -12\) maybe graphing isn't the best way to get a specific answer, just a nice ball park. Could you solve for x if I told you that the y terms are equal?

OpenStudy (anonymous):

how do i do that?

OpenStudy (fibonaccichick666):

So can you set the eq = to each other? We solved for y because we know that when these two lines intersect both their x and y coordinates are the same. If you set them =, you can find the x where the lines intersect.

OpenStudy (fibonaccichick666):

I have to go in 5 but I'll write a sample problem for you and post it in a min.

OpenStudy (fibonaccichick666):

http://www.educreations.com/lesson/view/jaylene2/9305862/ here you go a fully solved ex. problem

OpenStudy (fibonaccichick666):

good luck!

OpenStudy (anonymous):

Thank you for the help!